OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2. Find f(f(x)).

OpenStudy (anonymous):

Did you mean find f(g(x))?

OpenStudy (anonymous):

nopee, it says to find f(f(x))

OpenStudy (anonymous):

f(f(x)) = (x-2)-2 or f(f(x)) = x -4

OpenStudy (anonymous):

my answers are x-4, (x-2)-2, x^2-4, x^2-2..

OpenStudy (anonymous):

oh thanks!!!

OpenStudy (anonymous):

No problem.